OpenStudy (anonymous):

harry, peter, and 9 girls sit in a row. harry cant sit in an end seat. In how many ways can they be arranged?

OpenStudy (gabylovesyou):

do u have choices??

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

@satellite73 can U please help me its URGENT!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

can u help me with one question gabby? its a different one: how many 4 digit number can be formed tht are greater than 3000, and whos last digit is either a 3 or a 6

OpenStudy (anonymous):

about the first one. 11 people, harry can't sit in the first or last seat. so why don't we just say he can't sit in the first or second and make it easier 10 choice for first seat (not harry) then 9 for second (still not harry) then 2 people are seated leaving 9! ways for the rest of the chairs. make sense?

OpenStudy (anonymous):

in other words i think it is \[10\times 9\times 9!\]

OpenStudy (anonymous):

second on is easier. 7 digits for the thousands place 3,4,5,6,7,8,9 then 10 for the hundreds 10 for the tens and 2 for the ones place \[7\times 10\times 10\times 2\] i think is right

OpenStudy (anonymous):

okay thnks alot , but do u mind if i ask a couple more, cause im doing a timed test online, and really need help.

OpenStudy (anonymous):

oh and for the first question is it 10! times 9! times 9!

OpenStudy (anonymous):

no the first on is \[10\times 9\times 9!\] i think

OpenStudy (anonymous):

looks like a pretty damn big number http://www.wolframalpha.com/input/?i=10*9*9!

OpenStudy (anonymous):

second one is easier, it is 1400 unless i made a mistake no guarantees for these

OpenStudy (anonymous):

actually am fairly confident my first answer is right

OpenStudy (anonymous):

okay how abt this question: theres only six seats in an suv, one of the seats is modified to only fit a baby, and the other one is the drivers seat. the number of ways that 5 adults and a baby can be legally seated if only 4 of the adults are licensed is.

OpenStudy (anonymous):

4 for the drivers seat, then 4 times 3 times 2 for the remaining 4 seats forget the baby, only one choice there

OpenStudy (anonymous):

so i get \[4\times 4\times 3\times 2=96\]

OpenStudy (anonymous):

okay you are really help ful for tht u get a medal, and if u help me with a couple more please. i beg you. lol. i know i am being abit annoying but its all tht i get it in the end

OpenStudy (anonymous):

(n-3)P 2 = 110. what is the value of n

OpenStudy (anonymous):

is this supposed to be \(n-3\) permute 2 =110?

OpenStudy (anonymous):

this problem is straight forward. what two adjacent integers multiply to give 110? must be \[11\times 10\] so \(n-3=11\) and therefore \(n=14\)

OpenStudy (anonymous):

is this some on line course?

OpenStudy (anonymous):

class work

OpenStudy (anonymous):

its a practice test for marks.

OpenStudy (anonymous):

the final score in a soccer game is 6 to4 how many half time scores are possible?

OpenStudy (anonymous):

? anyone

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i guess we can count them 6 4, 6 3, 6 2, 61, 6 0 so 5 with the winning team at 6 5 with winning team at 5 etc etc \[7\times 5\]

OpenStudy (anonymous):

so 35

OpenStudy (anonymous):

@satellite73 thankyou so much for the help really appreciate it.

OpenStudy (anonymous):

yw, glad to help